When aqueous solutions of Na2SO4 and Pb(NO3)2 are mixed, PbSO4 precipitates. Calculate the limiting reagent and mass of PbSO4 when a 1.25 L solution of 0.0500 M Pb(NO3)2 and a 2.00 L solution of 0.0250 M Na2SO4 are mixed. (Hint: start first by balancing the equation of the reaction between Na2SO4 and Pb(NO3)2


1
Expert's answer
2022-04-15T10:05:04-0400

Na2SO4 + Pb(NO3)2 = 2NaNO3 + PbSO4


n(Na2SO4) = C(Na2SO4) * V(Na2SO4) = 0.0250 M * 2.00 L = 0.05 moles

n(Pb(NO3)2) = C(Pb(NO3)2) * V(Pb(NO3)2) =  0.0500 M * 1.25 L = 0.0625 moles


Pb(NO3)2 excess reagent

Na2SOlimiting reagent


m(PbSO4) = (m(Na2SO4)*Mr(PbSO4)) / Mr(Na2SO4) = n(Na2SO4)*Mr(PbSO4)

m(PbSO4) = 0.05 moles * 303 g/mol = 15.15 g